TMS Meeting Minutes, January 28, 2021

Attendees: Andrew, Breann, Heejoo, John, Matthieu, Trenton, Olga, Yang (Zoom Meeting)

Laser Alarm
  • The laser monitor email alarm is turned on and sending emails to the TMS group
  • Currently, the warning email appears to have a lag of 20 minutes
    • We would really like to have a delay time down to only one minute
    • Andrew wonders about the overhead involved in sampling more frequently
    • Currently, the iMonit system can only report at 10 minute intervals max (and many other systems on LBT already work with iMonit)
    • Heejoo suggests that he might be able to update this system with a microprocessor (e.g. Arduino) to speed up reports
  • At present, one email when the laser turns on and one email when the laser turns off
Andrew's Data Acquisition with Steve
  • Steve was instrumental in aligning the ch. 8 collimator
  • Unfortunately, ch. 11 is a problem, and we have excluded alignment as the source of the problem
    • The measurement laser doesn't appear to be transmitting, only the alignment laser
    • Heejoo notes that similar behavior exhibited by the MPO cassette (visible laser works fine, IR laser no go)
    • Andrew has removed ch. 11 from the data analysis
    • Steve found that ch. 11 was aligned to the wrong retro (-x instead of +y), has since been fixed
  • W/ 8 ch., Andrew took measurements with Etalon and the Etalon-calculated point positions, determined RxRyRzXYZ
  • Now working to compare these Etalon pose results to different channel combinations with the TCP/IP measurements
  • According to Joel, we shouldn't see any channel dropping; however, we see channel dropping at TMS, and Andrew is digging into this discrepancy
  • Andrew proposes that we should work to go on-sky with a stable system using the standard deviation channel test
  • When making a reference file after FPIA, need to be able to use the available channels
  • Once the reference is made, until making a new reference, we simply continue using this reference, and not using data if the channel drops
  • This should mitigate the risk of sending anything crazy
  • At the next time we run FPIA, we can choose then to drop a continuously dropped channel in the newly made reference
  • Interesting "clunk" behavior in X shown when moving very near to Zenith, though the magnitude is only a few microns (much smaller than anticipated)
    • See some hysteresis
    • By stepping in half degree steps, there is noticeable settling time
    • The settling took some minutes
    • All of the noted settling is only from flexure, no collimation was run
    • Could be happening because of rigid body motion of optics, or the collimator posts moving; however, for them to conspire to only affect X, this is unlikely
  • Didn't drop a single channel all night
  • Zenith, down in 5 degree steps down to 30 degrees, then back up
    • No gross hysteresis noted
  • Zenith to 30 to Zenith
    • Noted hystersis in X
    • Settling on the order of 40 microns over 3 minutes
  • Different hysteresis behavior depending on whether we are increasing in elevation or decreasing
  • Andrew suggests that we might use the guider cam or other to take shots as the telescope is moving and shortly after halting to measure the hysteresis with a bright star
  • Typical damping for a steel structure with bolts is 2%; we are seeing something different
  • Consider generating a pointing model with TMS for a system which is smoothly correcting the optics
    • This methodology could provide a quality control check of TMS

  • There are a couple of potential engineering nights at the start of February, and towards the second week
  • Would be very interesting to design a test to observe the hysteresis Andrew presented today
  • Could track on a star, moving in from one direction, then walk away and move back from another direction to test this
  • Once Yang's software is ready, Andrew wants to use closed-dome time to run overall system tests with TMS correcting to ensure the system manages all of the potential use cases well
  • Try to simulate all of the potential failure modes in closed-dome
  • Andrew would like to take another set of data soon (maybe tomorrow night) to measure hysteresis with the collimation model included

Fiber Collimators
  • Trenton and Heejoo currently working on a collimator study of off-the-shelf fiber collimators to determine a suitable solution
  • Andrew thinking of working with KiwiStar on developing a new collimator designed specifically for our purposes at a competitve price

  • Heejoo working on borrowing laser tracker from OEFF/Mirror Lab for testing of solid retro

-- TrentonBrendel - 28 Jan 2021
Topic revision: r1 - 28 Jan 2021, TrentonBrendel
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback